So my equation is 2/3x-22=14 I am doing what it tells me to do what ever it equals is what x = so that means it is 14 so I filled it in 22-2/3+14 solve
Answer by Alan3354(69443)

2/3x-22=14
Assuming you mean (2/3)x-22=14 and not 2/(3x) - 22=14:
2/3x-22=14
Multiply thru by 3
2x - 66 = 42
Add 66
2x = 108
x = 54
-------------
Less talk and more effort might be a good idea.